# Exploring the Energy Landscape of RBMs: Reciprocal Space Insights into Bosons, Hierarchical Learning and Symmetry Breaking

### v1

@@ 1 Introduction
Generative models are ubiquitous as they have emerged as powerful tools across multiple domains. The ongoing sprint to better models has led to a plethora of frameworks. In data-driven contexts, these models are designed to learn and replicate the underlying probability distributions of complex datasets. In particular, they have found significant applications in condensed matter and nuclear physics, where neural network-based ansätze are employed to model phase transitions and approximate ground-state wavefunctions[1, 2, 3, 4, 5, 6]. Similarly, generative models are employed in high-energy physics to address multiple problems, e.g., shower generation, jet generation and unfolding [7, 8, 9, 10]. Although generative AI techniques have achieved notable success in natural language processing and computer vision, our theoretical understanding of their learning dynamics and performance remains incomplete. Critical questions, such as the precise mechanisms through which these models capture data distributions, the robustness and fidelity of their approximations, and the criteria that determine the suitability of specific frameworks for particular datasets, remain largely unresolved. Addressing these issues is not merely an academic exercise; rather, it is essential for developing reliable and predictive generative models for complex physical systems.
@@
Diffusion models were first introduced as a type of hierarchical VAE, trained by optimizing a variational bound, similar to a VAE regularizer [15]. However, diffusion models outperform most frameworks in most tasks. Once the diffusion model is trained, the sampling process consists of first generating a Gaussian distributed random vector and then denoising it via the reverse diffusion process. This process is similar in spirit to the sampling process in an RBM.
RBMs are universal approximators of discrete distributions [16] and rely on computationally intensive Monte Carlo Markov Chain (MCMC) methods for training and sampling. By introducing a reciprocal space formulation for RBMs, we reveal a connection between RBMs, diffusion processes, and systems of coupled Bosons. Our analysis shows that at initialization, the RBM operates at a saddle point, where the local curvature is determined by the singular values of the weight matrix, whose distribution follows the Marc̆enko-Pastur law and exhibits rotational symmetry. During training, this rotational symmetry is broken due to hierarchical learning, where different degrees of freedom progressively capture features at multiple levels of abstraction. This leads to a symmetry breaking in the energy landscape, reminiscent of Landau’s theory. This symmetry breaking in the energy landscape is characterized by the singular values and the weight matrix eigenvector matrix. We derive the corresponding free energy in a mean-field approximation. We show that in the limit of infinite size RBM, the reciprocal variables are Gaussian distributed. Our findings indicate that in this regime, there will be some modes for which the diffusion process will not converge to the Boltzmann distribution. We demonstrate these phenomena with an RBM and MNIST and further discuss how various initialization strategies can influence training dynamics. Our findings not only bridge the gap between disparate generative frameworks but also shed light on the fundamental processes underpinning learning in deep generative models.
@@ 2 A brief introduction to Restricted Boltzmann Machines
All the numerical results regarding trained RBMs presented in this paper were obtained by training on MNIST dataset and are available in our Github repository [22]. The RBM training was done using the publicly available Julia code, as described in Ref. [18]. We trained five replicas for each RBM with hidden layer size $500,784,1200$ and $3000$. In Fig. 1 left panel we show the log-likelihood vs epochs to verify the maximization of the log-likelihood. We estimated the partition function using annealed importance sampling and reverse annealed importance sampling, which correspond to an upper and lower bound of the partition function [23, 24].
Figure 1: Left panel) Log-likelihood vs epochs for RBMs with hidden layer sizes $M=500,784,1200,3000$ The partition function was estimated using annealed importance sampling (AIS) and reverse AIS [23, 24]. Each data point corresponds to the average over five replicas and the error bars correspond to the standard deviation. Right panel) Image of number $3$ generated from a trained RBM and, via the transformations in Eq. (17a), projected onto the energy landscapes (clockwise starting at the upper left corner) $1,2,3,4,20,30,50,100,200,300,498$ and $500$. The magenta star marks the saddle point whereas the black pentagon corresponds to the image projection to reciprocal space.
@@ 3 Restricted Boltzmann machine in reciprocal space
Updating the weight matrix and the self-fields via Eqs. (13c) ultimately reduces the energy associated to the dataset points via the energy function (see Eq. (3)). The energy landscape described by the energy function has at most as many saddle points as there are units in the smaller partition. At initialization, the dataset points cluster around these saddle points; however, as training progresses, they deviate from the saddle points but never drift away, due to implicit constraint imposed by the binary nature of the variables. To show this, we project the RBM onto reciprocal space using singular value decomposition (SVD). By studying the eigenvector matrices and singular values of the weight matrix, we learn about the weight matrix properties and provide a physical interpretation of the singular values in the context of RBMs.
In general, the coupling matrix $\bm{W}$ is a rectangular, non-symmetric random matrix with Gaussian entries and standard deviation $\sigma$. We perform SVD on $\bm{W}$, such that $\bm{W}=\bm{U}\bm{\Sigma}\bm{V}^{t}=\sum_{\alpha}\lambda_{\alpha}|v_{\alpha}\rangle\langle h_{\alpha}|$; where $\bm{U}$ is a matrix composed of the eigenvectors of $\bm{W}\cdot\bm{W}^{t}$, i.e., $\bm{U}=\sum_{s}|v_{s}\rangle\langle c_{s}|$, where $\{|c_{s}\rangle\}$ denotes the canonical basis; $\Sigma$ is a rectangular matrix of size $N\times M$ containing the singular values $\{\lambda_{i}\}_{i=1}^{min(N,M)}$ on the diagonal; and $\bm{V}^{t}$ is a matrix composed of the eigenvectors of $\bm{W}^{t}\cdot\bm{W}$, i.e., $\bm{V}^{t}=\sum_{s}|c_{s}\rangle\langle h_{s}|$. Note that due to the orthogonality of $\bm{U}$ and $\bm{V}$, we have $\bm{U}\cdot\bm{U}^{t}=\bm{V}\cdot\bm{V}^{t}=\bm{I}$. Moreover, the eigenvector matrices $\bm{U}$ and $\bm{V}$ are Haar distributed, i.e., they are uniformly distributed over the group $O(N)$ and each eigenvector is uniformly distributed on the unit hypersphere [25]. The training process breaks this symmetry, as we will show later.
@@
$E_{i}(x_{i},y_{i})=\begin{cases}-a_{0i}x_{i}-b_{0i}y_{i}-\lambda_{i}x_{i}y_{i},\quad 1\leq i\leq min(N,M)\\ -a_{0i}x_{i},\quad min(N,M)<i\leq max(N,M)\\ \end{cases}$ (16)
@@
It is instructive to consider a dynamical system described by the Hamiltonian $E_{i}(x_{i},y_{i})$ [26]. In this scenario, the conjugate momenta are $p_{x_{i}}=a_{0i}+\lambda_{i}y_{i}$ and $p_{y_{i}}=b_{0i}+\lambda_{i}x_{i}$, respectively. Setting these momenta to zero yields the fixed point of the system. From the Hessian matrix we find that the eigenvalues correspond to ±$\pm\lambda_{i}$, indicating that the fixed point is a saddle point. Each of the $E_{i(\leq min(N,M))}$ corresponds to a hyperbolic paraboloid with saddle point at $(x_{0i},y_{0i})=(-b_{0i}/\lambda_{i},-a_{0i}/\lambda_{i})$ where the energy is equal to $b_{0i}a_{0i}/\lambda_{i}$. In other words, for each energy landscape $E_{i(\leq min(N,M))}$ the phase curves are such that there is always a direction where the energy decreases and tends to $\infty$- ∞.
@@
We further introduce a new set of variables $\{u_{i}\}_{i=1}^{\text{min(N,M)}},\{w_{i}\}_{i=1}^{\text{min(N,M)}}$ which relate to $x_{i},y_{i}$ for $1\leq i\leq min(N,M)$ via a translation to the saddle points and a rotation, namely,
@@
We can now give a physical interpretation of the weight matrix singular values. The singular values can be interpreted as the vibrational modes, since the curvature of each mode per energy landscape is ±$\pm\lambda$. In addition to modulating the stiffness of the mode, the singular values modulate the saddle point energy and its location. In Fig. 2 we show four energy landscapes in the $xy$-plane corresponding to the singular value indices $1,2,3$ and $498$ in the case of a trained RBM, and we also project the test set onto these landscapes. We do this for an untrained and a trained RBM. In the former, the dataset points cluster on the saddle point, whereas for the trained RBM this is not always the case although the points do not drift away from the saddle point.
@@
In practice, several empirical guidelines are used for initializing RBMs, many of which are supported by experimental evidence. For example, selecting the number of nodes in the hidden layer remains largely heuristic. A common recommendation is to limit the hidden layer to no more nodes than the visible layer, as exceeding this threshold may slow down training due to the increase in the number of operations per matrix multiplication. An additional reason to not have the hidden layer be much greater the the visible nodes is related to the singular values distribution. As we have shown previously, the singular values influence the stability of the energy landscape. The singular value spectrum is characterized by the Marc̆enko–Pastur law [25]:
@@
with $\lambda_{\pm}=\sqrt{\text{max}(N,M)\sigma^{2}}\pm\sqrt{q}$ and $q=\text{min}(N,M)\sigma^{2}$. In Fig. 3 we plot Eq. (20) for various RBM sizes. One can demonstrate that the maximum of $\varrho(\lambda)$ occurs at $\lambda_{max}=\sqrt{\lambda_{+}\lambda_{-}}$, implying $\lambda_{max}=\sigma\sqrt{|N-M|}$. This distribution increases with the difference between visible and hidden nodes as well as with increasing standard deviation $\sigma$. Consequently, a significant imbalance between the numbers of visible and hidden nodes and/or a broad range of coupling values accentuates the singular value distribution, thereby impacting the stability of the energy landscapes at initialization.
@@ 4 Partition function in reciprocal space
We first construct the partition function of the reciprocal space in terms of the variables $\{u_{i}\}_{i=1}^{\text{min(N,M)}},\{w_{i}\}_{i=1}^{\text{min(N,M)}}$ and $\{x_{j}\}_{j=\text{min}(N,M)+1}^{\text{max}(N,M)}$. We follow a similar path to the replica method in the case of a single replica. Applying the replica method to RBMs has been done previously in Ref. [27].
@@
In this context, any binary state has an equal probability $2^{-N-M}$, which implies that $\langle v_{k}\rangle=\langle h_{k}\rangle=1/2$ for all $k=1,...,N+M$.
@@ 4.1 Fokker-Planck Equation and coupled Bosons
is the stationary solution of the Fokker-Planck Equation:
@@
We assume the over-damped regime with the friction coefficient set to $\gamma=1$ such that $D=1/\beta$. It is straightforward to show that Eq. (32) is the stationary solution to Eq. (33). By the variable separation method we propose the ansatz $P(\bm{z},t)=\phi(\bm{z})f(t)$, which substituting in Eq. (33) and dividing by the ansatz leads to:
@@
$\omega_{i}=\begin{cases}\frac{k_{ii}-\lambda\beta}{2}\quad\text{for }i=1,...,M\\ \frac{k_{ii}+\lambda\beta}{2}\quad\text{for }i=M+1,...,2M\\ \frac{k_{ii}}{2}\quad\text{for }i=2M+1,...,N\\ \end{cases}$ (40)
@@
$z_{i0}=\begin{cases}\frac{\mu_{z_{i}}k_{ii}}{2\omega_{i}}\quad\text{for }i=1,...,2M\\ \frac{\beta a_{i}}{2\omega_{i}}+\mu_{z_{i}}\quad\text{for }i=2M+1,...,N\\ \end{cases}$ (41)
@@ 5.1 Rotational Symmetry
RBMs are typically initialized by drawing each element of the weight matrix independently from a Gaussian distribution with mean 00 and standard deviation $0.01$ [29]. This initialization ensures that the RBM begins in a paramagnetic phase [30, 27]. Moreover, this procedure inherently confers rotational invariance to the joint probability density of the weight elements. In fact, the joint probability density of the weight elements is given by
@@
We probe the rotational symmetry in trained RBMs by applying random rotations on the eigenvector matrices $\bm{U}$ and $\bm{V}$, and studying the effect on the weight matrix distribution. We build a rotation matrix using the method described in Ref. [31]. This method performs rotations in $n-$dimensions around any arbitrary $(n-2)$-dimension subspace. We then apply $N*0.1$ and $M*0.1$ consecutive random rotations to random subspaces of $(N-2)$-dimension and $(M-2)$-dimension of $\bm{U}$ and $\bm{V}$, respectively. We then compare the weight matrix before and after the random rotations in the case of an initialized RBM with 500 hidden nodes, and trained RBMs with $500,784,1200$ and $3000$ hidden nodes by plotting the weight matrix PDF in Fig. 4. We also show the Jensen divergence between the rotated and non-rotated weight matrix. Each point corresponds to 5 replicas. The black dashed line corresponds to a non-trained RBM. This shows that the training process breaks the rotational symmetry.
@@
Figure 6: Gibbs sampled data. Same data after rotating by $\pi$ the eigenvector matrix $\bm{U}$ around all but the two dimensions associated with the largest singular values. Same data after 10 consecutive rotations around any $(N-2)$-dimension where the dimensions associated with the five largest singular values are always included and the remaining $N-7$ dimensions are randomly selected.

### v2

@@ 1 Introduction
Generative models are ubiquitous as they have emerged as powerful tools across multiple domains. The ongoing sprint to better models has led to a plethora of frameworks. In data-driven contexts, these models are designed to learn and replicate the underlying probability distributions of complex datasets. They have found significant applications in condensed matter [1], particularly, in quantum many-body [2, 3, 4], in thermodynamics [5] and nuclear physics [6], where neural network-based ansätze are employed to model phase transitions and approximate ground-state wavefunctions. Similarly, generative models are employed in high-energy physics to address multiple problems, e.g., calorimeter shower generation, jet generation [7, 8, 9] and unfolding [10, 11]. Although generative AI techniques have achieved notable success in natural language processing and computer vision, our theoretical understanding of their learning dynamics and performance remains incomplete. Critical questions, such as the precise mechanisms through which these models capture data distributions, the robustness and fidelity of their approximations, and the criteria that determine the suitability of specific frameworks for particular datasets, remain largely unresolved. Addressing these issues is not merely an academic exercise; rather, it is essential for developing reliable and predictive generative models for complex physical systems.
@@
Diffusion models were first introduced as a type of hierarchical VAE, trained by optimizing a variational bound, similar to a VAE regularizer [16]. However, diffusion models outperform most frameworks in most tasks. Once the diffusion model is trained, the sampling process consists of first generating a Gaussian distributed random vector and then denoising it via the reverse diffusion process [17]. This process is similar in spirit to the sampling process in an RBM , however, diffusion models in general violate detailed balance.
RBMs are universal approximators of discrete distributions [18] and rely on computationally intensive Monte Carlo Markov Chain (MCMC) methods for training and sampling. In recent times, RBMs have gained attention within different fields in physics, in particular, from the statistical physics community [19, 20], but also from quantum field theory [21]. Special interest is given to understanding the phase diagram of RBM learning [22, 23]. By introducing a reciprocal space formulation for RBMs, we reveal a connection between RBMs, diffusion processes, and systems of coupled Bosons. Our analysis shows that at initialization, the RBM operates at a saddle point, where the local curvature is determined by the singular values of the weight matrix, whose distribution follows the Marc̆enko-Pastur law and exhibits rotational symmetry. During training, this rotational symmetry is broken due to hierarchical learning, where different degrees of freedom progressively capture features at multiple levels of abstraction. This leads to a symmetry breaking in the energy landscape, reminiscent of Landau’s theory. This symmetry breaking in the energy landscape is characterized by the singular values and the weight matrix eigenvector matrix. We derive the corresponding free energy in a mean-field approximation. We show that in the limit of infinite size RBM, the reciprocal variables are Gaussian distributed. Our findings indicate that in this regime, there will be some modes for which the diffusion process will not converge to the Boltzmann distribution. We demonstrate these phenomena with an RBM and MNIST and further discuss how various initialization strategies can influence training dynamics. Our findings not only bridge the gap between disparate generative frameworks but also shed light on the fundamental processes underpinning learning in deep generative models. Additionally, recent work has demonstrated spontaneous breaking of reverse and permutation symmetries in the RBM weight‐matrix distribution at distinct critical ratios of dataset size to visible layer size [22, 20].
@@ 2 A brief introduction to Restricted Boltzmann Machines
All the numerical results regarding trained RBMs presented in this paper were obtained by training on MNIST dataset [29, 30] and are available in our Github repository [31]. The RBM training was done using the publicly available Julia code, as described in Ref. [25]. We trained five replicas for each RBM with hidden layer size $500,784,1200$ and $3000$. In Fig. 1 left panel we show the log-likelihood vs epochs to verify the maximization of the log-likelihood. We estimated the partition function using annealed importance sampling and reverse annealed importance sampling, which correspond to an upper and lower bound of the partition function [32, 33].
Figure 1: Left panel) Log-likelihood vs epochs for RBMs with hidden layer sizes $M=500,784,1200,3000$ . The partition function was estimated using annealed importance sampling (AIS) and reverse AIS [32, 33]. Each data point corresponds to the average over five replicas and the error bars correspond to the standard deviation. Right panel) Image of number $3$ generated from a trained RBM and, via the transformations in Eq. (17a), projected onto the energy landscapes (clockwise starting at the upper left corner) $1,2,3,4,20,30,50,100,200,300,498$ and $500$. The magenta star marks the saddle point whereas the black pentagon corresponds to the image projection to reciprocal space.
@@ 3 Restricted Boltzmann machine in reciprocal space
Updating the weight matrix and the self-fields via Eqs. (13c) ultimately reduces the energy associated to the dataset points via the energy function (see Eq. (3)). The energy landscape described by the energy function has at most as many saddle points as there are units in the smaller partition. At initialization, the dataset points cluster around these saddle points; however, as training progresses, they deviate from the saddle points but never drift away, due to implicit constraint imposed by the binary nature of the variables. To show this, we project the RBM onto reciprocal space by projecting the state vectors via the eigenvectors of weight matrix using singular value decomposition (SVD). Projecting the system onto this reciprocal space, we obtain a new set of variables with single pairwise interaction each. These new variables capture collective patterns weighed by the eigenvectors. By studying the eigenvector matrices and singular values of the weight matrix, we learn about the weight matrix properties and provide a physical interpretation of the singular values in the context of RBMs. We highlight that in Refs. [19, 34] a similar approach has been proposed as means to study the learning dynamics in RBMs. In particular, the authors study the learning dynamics in terms of the singular values of the weight matrix.
Upon initialization, the coupling matrix $\bm{W}$ is a rectangular, non-symmetric random matrix with Gaussian entries and standard deviation $\sigma$. We perform SVD on $\bm{W}$, such that $\bm{W}=\bm{U}\bm{\Sigma}\bm{V}^{t}=\sum_{\alpha}\lambda_{\alpha}|v_{\alpha}\rangle\langle h_{\alpha}|$; where $\bm{U}$ is a matrix composed of the eigenvectors of $\bm{W}\cdot\bm{W}^{t}$, i.e., $\bm{U}=\sum_{s}|v_{s}\rangle\langle c_{s}|$, where $\{|c_{s}\rangle\}$ denotes the canonical basis; $\Sigma$ is a rectangular matrix of size $N\times M$ containing the singular values $\{\lambda_{i}\}_{i=1}^{min(N,M)}$ on the diagonal; and $\bm{V}^{t}$ is a matrix composed of the eigenvectors of $\bm{W}^{t}\cdot\bm{W}$, i.e., $\bm{V}^{t}=\sum_{s}|c_{s}\rangle\langle h_{s}|$. Note that due to the orthogonality of $\bm{U}$ and $\bm{V}$, we have $\bm{U}\cdot\bm{U}^{t}=\bm{V}\cdot\bm{V}^{t}=\bm{I}$. Moreover, the eigenvector matrices $\bm{U}$ and $\bm{V}$ are Haar distributed, i.e., they are uniformly distributed over the group $O(N)$ and each eigenvector is uniformly distributed on the unit hypersphere [35]. The training process breaks this symmetry, as we will show later.
@@
[…]
@@
It is instructive to consider a dynamical system described by the potential $E_{i}(x_{i},y_{i})$, such that the state of the system is given by $x_{i}$ and $y_{i}$ and a velocity field $(\dot{x}_{i},\dot{y}_{i})=(\partial E_{i}(x_{i},y_{i})/\partial x_{i},\partial E_{i}(x_{i},y_{i})/\partial y_{i})$ [36]. The dynamics of the system are described by the set of equations:
$\begin{pmatrix}\dot{x}_{i}\\ \dot{y}_{i}\end{pmatrix}=-\begin{pmatrix}a_{0i}\\ b_{0i}\end{pmatrix}-\begin{pmatrix}0&\lambda\\ \lambda&0\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}$ (18)
Setting the velocity field to zero yields the fixed point of the system. From the Hessian matrix we find that the eigenvalues correspond to $\pm\lambda_{i}$, indicating that the fixed point is a saddle point. Each of the $E_{i(\leq min(N,M))}$ corresponds to a hyperbolic paraboloid with saddle point at $(x_{0i},y_{0i})=(-b_{0i}/\lambda_{i},-a_{0i}/\lambda_{i})$ where the energy is equal to $b_{0i}a_{0i}/\lambda_{i}$. In other words, for each energy landscape $E_{i(\leq min(N,M))}$ the phase curves are such that there is always a direction where the energy decreases and tends to -$\infty$. Another way to see this is by noticing that the energy landscape $E_{i}(x_{i},y_{i})$ is the equation of a conic section, in particular, of a hyperbola.
@@
We define a set of variables $\{u_{i}\}_{i=1}^{\text{min(N,M)}},\{w_{i}\}_{i=1}^{\text{min(N,M)}}$ which relate to $x_{i},y_{i}$ for $1\leq i\leq min(N,M)$ via a translation to the saddle points and a rotation, namely,
@@
We can now give a physical interpretation of the weight matrix singular values. The singular values can be interpreted as the vibrational modes, since the curvature of each mode per energy landscape is $\pm\lambda$. In addition to modulating the stiffness of the mode, the singular values modulate the saddle point energy and its location. In Fig. 2 we show four energy landscapes in the $xy$-plane corresponding to the singular value indices $1,2,3$ and $498$ in the case of a trained RBM, and we also project the test set onto these landscapes. We do this for an untrained and a trained RBM. In the former, the dataset points cluster on the saddle point, whereas for the trained RBM this is not always the case although the points do not drift away from the saddle point.
@@
In practice, several empirical guidelines are used for initializing RBMs, many of which are supported by experimental evidence. For example, selecting the number of nodes in the hidden layer remains largely heuristic. A common recommendation is to limit the hidden layer to no more nodes than the visible layer, as exceeding this threshold may slow down training due to the increase in the number of operations per matrix multiplication. An additional reason to not have the hidden layer be much greater the the visible nodes is related to the singular values distribution. As we have shown previously, the singular values influence the stability of the energy landscape. As noted previously in Refs. [19, 34], the singular value spectrum is characterized by the Marc̆enko–Pastur law [35]:
@@
with $\lambda_{\pm}=\sqrt{\text{max}(N,M)\sigma^{2}}\pm\sqrt{q}$ and $q=\text{min}(N,M)\sigma^{2}$. In Fig. 3 we plot Eq. (21) for various RBM sizes. One can demonstrate that the maximum of $\varrho(\lambda)$ occurs at $\lambda_{max}=\sqrt{\lambda_{+}\lambda_{-}}$, implying $\lambda_{max}=\sigma\sqrt{|N-M|}$. This distribution increases with the difference between visible and hidden nodes as well as with increasing standard deviation $\sigma$. Consequently, a significant imbalance between the numbers of visible and hidden nodes and/or a broad range of coupling values accentuates the singular value distribution, thereby impacting the stability of the energy landscapes at initialization. Note that if the weight matrix is initialized with $\sigma^{2}\sim 1/\text{max}(N,M)$, the extensive scaling with the dimensions drops out and only a dependence on $\text{max}(N,M)/\text{max}(N,M)\leq 1$ remains.
@@ 4 Partition function in reciprocal space
We first construct the partition function of the reciprocal space in terms of the variables $\{u_{i}\}_{i=1}^{\text{min(N,M)}},\{w_{i}\}_{i=1}^{\text{min(N,M)}}$ and $\{x_{j}\}_{j=\text{min}(N,M)+1}^{\text{max}(N,M)}$. We follow a similar path to the replica method in the case of a single replica. Applying the replica method to RBMs has been done previously in Refs. [34, 37].
@@
In this context, any binary state has an equal probability $2^{-N-M}$, which implies that $\langle v_{k}\rangle=\langle h_{k}\rangle=1/2$ for all k=1,,N+Mk=1,...,N+M.
@@ 4.1 Fokker-Planck Equation and coupled Bosons
In this sense, we now have mapped the RBM distribution to a Boltzmann distribution of a set of continuous variables described by the effective potential $U_{eff}(\bm{z})$ immersed in a heat bath with inverse temperature $\beta$. Therefore, Eq (33) is the stationary solution of the Fokker-Planck equation [38]:
@@
We assume the over-damped regime with the friction coefficient set to $\gamma=1$ such that the diffusion coefficient, defined via the fluctuation-dissipation theorem, becomes $D=1/\beta$. We also highlight that it may be possible to recast the training dynamics under a framework of Dyson Brownian dynamics as done in Ref. [39]. It is straightforward to show that Eq. (33) is the stationary solution to Eq. (34). By the variable separation method we propose the ansatz $P(\bm{z},t)=\phi(\bm{z})f(t)$, which substituting in Eq. (34) and dividing by the ansatz leads to:
@@
[…]
@@
[…]
@@ 5.1 Rotational Symmetry
RBMs are typically initialized by drawing each element of the weight matrix independently from a Gaussian distribution with mean 0 and standard deviation $0.01$ [41]. This initialization ensures that the RBM begins in a paramagnetic phase [42, 37]. Moreover, this procedure inherently confers rotational invariance to the joint probability density of the weight elements. In fact, the joint probability density of the weight elements is given by
@@
We probe the rotational symmetry in trained RBMs by applying random rotations on the eigenvector matrices $\bm{U}$ and $\bm{V}$, and studying the effect on the weight matrix distribution. We build a rotation matrix using the method described in Ref. [43]. This method performs rotations in $n$-dimensions around any arbitrary (n2)(n-2)-dimension subspace. We then apply N0.1N*0.1 and M0.1M*0.1 consecutive random rotations to random subspaces of (N2)(N-2)-dimension and (M2)(M-2)-dimension of $\bm{U}$ and $\bm{V}$, respectively. We then compare the weight matrix before and after the random rotations in the case of an initialized RBM with 500 hidden nodes, and trained RBMs with $500,784,1200$ and $3000$ hidden nodes by plotting the weight matrix PDF in Fig. 4. We also show the Jensen divergence between the rotated and non-rotated weight matrix. Each point corresponds to 5 replicas. The black dashed line corresponds to a non-trained RBM. This shows that the training process breaks the rotational symmetry. In addition to rotational symmetry, reverse symmetry likewise breaks during training. Moreover, it is easy to note that hierarchical learning is incompatible with permute symmetry, which ultimately breaks during training, as shown in [22, 20].
@@
Figure 6: Gibbs sampled data. Same data after rotating by $\pi$ the eigenvector matrix $\bm{U}$ around all but the two dimensions associated with the largest singular values. Same data after 10 consecutive rotations around any (N2)(N-2)-dimension where the dimensions associated with the five largest singular values are always included and the remaining N7N-7 dimensions are randomly selected.